OpenStudy (anonymous):

Use the Subtraction Property of Equality to complete the following statement: if 10x+6=21, then_____=15 10x + 6 10x – 6 10x 1.5

OpenStudy (vera_ewing):

B. 10x-6

OpenStudy (anonymous):

B. 10x-6

OpenStudy (superflytyguy):

10x

OpenStudy (superflytyguy):

Subtract 6 on both sides

OpenStudy (superflytyguy):

10x + 6 - 6 = 21 - 6 10x = 15

OpenStudy (anonymous):

so the answer is 10x-6??